OpenStudy (anonymous):

how do you factor 4x^-9x-9=0

hero (hero):

Did you mean 4x^2 - 9x - 9 = 0

OpenStudy (anonymous):

yes i do.

hero (hero):

In general, find two numbers that multiply to get a*c and add to get b In this case, Find two numbers that multiply to get 4*9 and add to get -9 Those two numbers are -12 and 3 Now split the middle term using those two numbers, then continue solving for x: 4x^2 - 12x + 3x -9 = 0 4x(x-3)+3(x-3) = 0 (x-3)(4x+3) = 0 x = 3, -3/4

OpenStudy (anonymous):

omg thank you! you are a life saver hero. :)